OpenStudy (anonymous):

Determine whether equation has the given ordered pair as a solution. 3q-2p=-8; (1,-2) please show work!!

OpenStudy (anonymous):

substitute p = 1 and q = -2 and see whether LHS = RHS

OpenStudy (anonymous):

i wanted to know was this corrected 3x-2 -2x-1 3(-2)-2(-1) -6-2- -8

OpenStudy (anonymous):

-6-2=-8

OpenStudy (anonymous):

actually, if p=1 and q=-2 so, we are going to substitute p=1 and q=-2 in the equation 3q-2p=-8 3(-2)-2(1)=-8 -6-2=-8 -8=-8